OpenStudy (anonymous):

can you help find the area of a circle centered at (1,1) that passes through the point (-2,6).

OpenStudy (jdoe0001):

the circle's center is at (1,1) and passes through (-2,6) meaning that the distance between (1,1) and (-2,6) is the "radius" recall that the area of a circle = \(\bf \pi r^2\) so find the distance for those 2 points to get the radius, then use the area of a circle's formula \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ 1}})\quad &({\color{red}{ -2}}\quad ,&{\color{blue}{ 6}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (anonymous):

 find r by distance formula,then area \[A=\pi r^2\]